# A Two-Stage GPU Kernel Tuner Combining Semantic Refactoring and Search-Based Optimization

### v1

@@
Abstract:  GPU code optimization is a key performance bottleneck for HPC workloads as well as large-model training and inference. Although compiler optimizations and hand-written kernels can partially alleviate this issue, achieving near–hardware-limit performance still relies heavily on manual code refactoring and parameter tuning. Recent progress in LLM-agent-based kernel generation and optimization has been reported, yet many approaches primarily focus on direct code rewriting, where parameter choices are often implicit and hard to control, or require human intervention, leading to unstable performance gains. This paper introduces a template-based rewriting layer on top of an agent-driven iterative loop: kernels are semantically refactored into explicitly parameterizable templates, and template parameters are then optimized via search-based autotuning, yielding more stable and higher-quality speedups. Experiments on a set of real-world kernels demonstrate speedups exceeding 3× in the best case. We extract representative CUDA kernels from SGLang as evaluation targets; the proposed agentic tuner iteratively performs templating, testing, analysis, and planning, and leverages profiling feedback to execute constrained parameter search under hardware resource limits. Compared to agent-only direct rewriting, the template-plus-search design significantly reduces the randomness of iterative optimization, making the process more interpretable and enabling a more systematic approach toward high-performance configurations. The proposed method can be further extended to OpenCL, HIP, and other backends to deliver automated performance optimization for real production workloads.
Keywords: search-based optimization, GPU code tuning, multi-agent systems
@@ 1 Introduction
As AI workloads rapidly grow in scale and diversify in form, GPU kernel performance has become a critical bottleneck for the efficiency of training and inference systems. Although large language models have shown strong capability in code generation and program rewriting, reliably producing GPU kernels that compile, are correct, and approach expert-level performance remains challenging. The root cause is a highly combinatorial optimization space in which structural transformations are tightly coupled with hardware resource constraints. Free-form rewriting directly at the source level is difficult to control and reproduce, can easily introduce correctness deviations, and often leads to volatile performance. This paper presents a two-level GPU code tuning approach that combines semantic refactoring and search-based optimization. We first use measurement- and constraint-driven refactoring to transform a kernel into an explicitly parameterizable template representation, so optimization intent is systematized and managed through an explicit parameter space. We then perform search-based fine-grained autotuning within the hardware-feasible region, with a closed feedback loop consisting of compilation and execution, correctness verification, and performance evaluation. The resulting feedback further drives semantic-level rewriting to continuously raise the structural performance ceiling, enabling stable convergence toward high-performance implementations while preserving correctness.
Our contributions are as follows.
1. We propose an automated GPU kernel optimization framework that integrates hardware resource awareness, template-based refactoring, and search-based fine-grained autotuning, turning the optimization process from unconstrained rewriting into reproducible parameter search.
2. We introduce a layered optimization mechanism in which semantic-level rewriting improves the structural upper bound, while fine-grained parameter tuning approaches the best mapping, achieving stable performance gains on the target hardware while maintaining correctness.
3. We build an end-to-end closed-loop workflow with multi-agent collaboration that unifies semantic-level refactoring, templating, tuning search, correctness checking, and performance measurement, and uses measurement feedback to drive iterative semantic-level rewriting.
@@ 2 Related Work
Multi-Agent Systems (MAS). MAS consists of multiple interacting agents that collaborate to solve complex shared tasks beyond the capability of a single agent. This paradigm naturally aligns with programming-oriented workflows: complex software development processes can be decomposed into subtasks such as planning, implementation, testing, and profiling, which can then be assigned to different agents for coordinated and iterative progress. In recent years, a growing body of work has explored multi-agent frameworks (e.g., AutoGen, Trace, MetaGPT), demonstrating strong capability and empirical effectiveness on benchmark tasks including mathematical reasoning and code generation[ref2, ref3, ref4, ref5, ref6, ref7, ref8, ref9, ref10]. However, systematic application of MAS to parallel code optimization remains relatively limited. A fundamental reason is that performance optimization for modern accelerators relies heavily on specialized knowledge of hardware microarchitecture, parallel execution models, and memory hierarchies, with complex optimization objectives and constraints. As a result, while multi-agent approaches are promising, they also introduce new challenges in modeling and coordination.
@@
Compiler- and learning-based methods for GPU kernel optimization. For decades, parallel code optimization has been driven primarily by compiler frameworks and domain-specific languages (DSLs). Representative systems include Halide, TVM, MLIR, TensorFlow XLA, and NVIDIA CUTLASS[ref11, ref12, ref13, ref14, ref15, ref16, ref17, ref18, ref19]. These systems provide high-level abstractions for expressing tensor computations and leverage compiler optimization pipelines to perform scheduling transformations, operator fusion, and memory-hierarchy mapping, thereby reducing the cost of manual optimization.
@@
To further approach expert-level performance, autotuning has become a key technique. Frameworks such as AutoTVM, Ansor, AMOS, KernelTuner, and OpenTuner explore large optimization spaces via search and machine learning to identify hardware-specialized configurations[ref20, ref21, ref22, ref23, ref24]. Building on this foundation, recent systems (e.g., Triton, Mirage, ThunderKittens) further expand the design space of the “DSL + tuning” paradigm. For instance, Triton introduces a tile-level intermediate representation coupled with autotuning, enabling generated kernels to reach performance close to hand-optimized implementations under certain conditions[ref25, ref26, ref27]. Despite this progress, compiler/DSL-driven approaches still face two prominent limitations. First, without substantial tuning effort, they often struggle to consistently achieve expert-level performance, and their generalization across hardware platforms remains insufficient[ref27]. Second, these systems typically rely on relatively rigid compilation pipelines, which incur high development and maintenance costs and demand significant engineering investment. In other words, there remains a clear tension between high-performance attainability and system scalability/portability.
@@
LLM-driven high-performance code generation and optimization. Early work on large language model (LLM)-based code generation primarily targeted general-purpose programming tasks (e.g., AlphaCode) and demonstrated promising potential[ref28]. In recent years, the focus has increasingly shifted toward performance-oriented, domain-specific code generation, spanning vectorization[ref29, ref30], assembly-level optimization[ref31], parallel program generation for DSLs[ref32, ref9, ref33], and tensor program optimization[ref34]. Among these directions, automatic generation of high-performance GPU kernels has emerged as a particularly active line of research[ref35, ref36]. Unlike general-purpose code generation, performance optimization offers verifiable reward signals (e.g., latency, throughput, and bandwidth utilization). Consequently, an iterative optimization paradigm of “generate–evaluate–improve” has become increasingly prevalent: the model first produces candidate kernels, which are then refined through a feedback loop involving compilation checks, correctness verification, runtime profiling, or self-reflection[ref1, ref37, ref9, ref38, ref39]. However, existing LLM- or multi-agent-based tuning processes largely rely on LLM-driven feedback loops and lack an interpretable and controllable mechanism for fine-grained parameter tuning. This often causes key parameter boundaries and interaction effects under hardware resource constraints to be missed, leading to incomplete exploration and high variance in performance.
@@ 3.1 Task Definition
The task considered in this study is as follows. Given an original GPU kernel $S$ (CUDA/OpenCL/HIP) and a target hardware device $h$, we first perform template-based refactoring on $S$, explicitly parameterizing its optimizable dimensions into a searchable fine-grained parameter vector $\theta$, which yields a family of parameterized implementations $\{S_{\theta}\}$. We then conduct search-based autotuning over $\theta$ under hardware resource feasibility constraints to obtain candidate optimal implementations, followed by correctness verification and performance measurement. Finally, guided by the measured data, we derive semantic-level optimization suggestions targeting hardware resource utilization, and apply semantics-preserving structural rewriting to the original implementation to improve the starting point and the achievable performance ceiling for the next tuning round. This procedure can be executed iteratively, and ultimately outputs a performance-optimal kernel $S^{\prime}$ while preserving functional correctness.
Template-based refactoring and fine-grained parameter tuning. Let the input domain be $X$ and the output space be $Y$. The original kernel is:
$S:X\to Y.$
Starting from an implementation of $S$, we perform template-based refactoring via non-semantic rewriting and introduce a quantifiable, searchable fine-grained parameter vector $\theta$, rewriting it into a parameterized family $\{S_{\theta}\}$. Formally, let the templateization and instantiation process be:
$S_{\theta}=\mathrm{Instantiate}\bigl(\mathrm{Templateize}(S,\theta)\bigr),\quad\dots\quad S_{\theta}:X\to Y.$
Here, $\theta$ captures hardware mapping and micro-tuning parameters such as block/work-group shape, vector width, unrolling factors, shared/local memory usage, and pipeline staging.
3.2 Hardware Resources and Tuning Constraints
Hardware resource feasibility. Let the target hardware be $h$, whose resource upper-bound vector is $r(h)$, including limits on registers/shared memory, the maximum number of threads or work-group size, vector alignment constraints, and so on. The resource consumption of a template instance on hardware $h$ is denoted by $u(\theta;h)$, and we require:
$u(\theta;h)\preceq r(h).$
This constraint captures how hardware resource awareness shrinks the search space and ensures that candidate implementations are compilable and executable.
Tuning. Within the resource-feasible region, we perform search-based autotuning over $\theta$ and obtain a candidate optimal parameter setting:
$\theta^{*}\in\arg\min_{\theta:\,u(\theta;h)\preceq r(h)}T_{h}(S_{\theta},x),$
where $T_{h}(S_{\theta},x)$ denotes the runtime of kernel $S_{\theta}$ on hardware $h$ for input $x$. The corresponding candidate output kernel is:
$S^{\prime}=S_{\theta^{*}}.$
Correctness. Ideally, the templateized-and-instantiated kernel should be equivalent to the original kernel over the entire input domain:
$\forall x\in X:\,S^{\prime}(x)\equiv S(x).$
To account for floating-point error, we allow approximate equivalence under a metric $d$ and tolerance $\varepsilon\geq 0$:
$\forall x\in X:\,d(S^{\prime}(x),S(x))\leq\varepsilon.$
In practice, correctness of $S^{\prime}$ is validated using a finite test set $T=\{x_{i}\}_{i=1}^{m}$:
$\max_{1\leq i\leq m}d(S^{\prime}(x_{i}),S(x_{i}))\leq\varepsilon.$
Performance measurement and evaluation. Let $T_{h}(S,x)$ denote the runtime of kernel $S$ on hardware $h$ for input $x$. For each $x\in T$, we focus on the best performance over the feasible parameter set:
$T_{h}^{*}=\min_{\theta\in\text{feasible}}T_{h}(S_{\theta},x),$
and define the speedup relative to the original kernel $S$ as:
$\sigma^{*}(x)=\frac{T_{h}(S,x)}{T_{h}^{*}(x)}=\max_{\theta\in\text{feasible}}\frac{T_{h}(S,x)}{T_{h}(S_{\theta},x)}.$
We summarize speedups over the test set $T$ using the geometric mean:
@@ 3.2 Hardware Resources and Tuning Constraints
Here, $\theta\cdot\text{feasible}$” denotes parameter assignments that simultaneously satisfy correctness and resource constraints, i.e.,
$\theta\cdot\text{feasible}\iff d(S_{\theta},S(x))\leq\varepsilon\land u(\theta;h)\preceq r(h).$
Semantic-level rewriting feedback. After correctness verification and performance measurement, we aggregate measurement results to produce semantic-level optimization suggestions oriented toward hardware resource utilization. These suggestions then guide semantics-preserving structural rewriting of the original implementation to improve resource utilization and raise the achievable performance ceiling.
For each input $x$, we collect performance measurements such as runtime, bandwidth, occupancy, and memory access efficiency. Let the measurement procedure be $\boldsymbol{M}$, and the measurement result be:
$m(x)=\boldsymbol{M}(S,x;h).$
Based on the measured data $m(x)$, we generate high-level semantic optimization suggestions and use them to guide rewriting decisions. Let the suggestion generator be $\boldsymbol{A}$ and the rewriting operator be $\boldsymbol{R}$:
$g=\boldsymbol{A}\bigl(m(\cdot)\bigr),\quad S^{hi}=\boldsymbol{R}(S,g).$
Here, $g$ denotes semantic-level suggestions targeting hardware resource utilization, such as increasing parallelism and occupancy, enhancing data reuse and locality, reducing global memory traffic and branch divergence, improving memory coalescing and alignment, and adjusting the compute-to-memory balance. The rewritten implementation $S^{hi}$ is used as the input to the next round of template-based refactoring and fine-grained parameter tuning.
3.3 Multi-agent implementation

### v2

@@
Abstract:  To address these challenges, we propose a two-level GPU code tuner that combines semantic refactoring with search-based optimization. Building on an agentic iterative loop, our method introduces template-based rewriting that explicitly refactors candidate kernels into parameterizable templates, and then performs search-based auto-tuning over the template parameters to obtain more stable and higher-quality performance improvements. We extract real-world kernels from SGLang as evaluation targets. The proposed agentic tuner completes templating, testing, analysis, and planning within the iterative loop, and conducts parameter search under hardware constraints guided by profiling signals. Experiments show that, compared with SGLang’s native kernel implementations, our method achieves over 3× speedup on a set of real kernels. Compared with agent-only direct rewriting, templating plus search significantly reduces iteration randomness, makes the optimization process more interpretable, and approaches high-performance configurations in a more systematic manner. The approach can be further extended to backends such as OpenCL and HIP, providing automated performance optimization for real serving workloads.
Keywords: search-based optimization, CUDA, GPU, auto-tuning, agents
@@ 1 Introduction
As AI workloads rapidly evolve in scale and structure, GPU kernel performance has increasingly become a critical bottleneck for training and inference system efficiency. Although large language models have demonstrated strong capabilities in code generation and program rewriting, consistently producing GPU kernels that are compilable, correct, and close to expert-level performance remains challenging. The root cause lies in the highly combinatorial optimization space and the tight coupling between structural transformations and hardware resource constraints. Free-form source-level rewriting is therefore difficult to control and reproduce, and can easily introduce semantic deviations and performance variability.
To address this issue, we leverage the automation capability of multi-agent iterative optimization and develop a unified formulation that integrates semantics-preserving refactoring with resource-feasible search. Based on this formulation, we design a two-level GPU code tuner that combines semantic refactoring and parameter search. Specifically, guided by measurement feedback and hardware constraints, we first refactor candidate kernels into explicit parameterizable templates, so that optimization intent is mapped into a manageable and reproducible parameter space. We then perform search-based fine-grained tuning within the resource-feasible region, closing the loop through compilation/execution, correctness verification, and performance evaluation. Furthermore, we use measurement feedback to drive semantics-level structural rewriting, continuously raising the reachable performance upper bound and thereby stably approaching high-performance implementations while preserving correctness.
Our contributions are as follows:
(1) We build an automated optimization framework for GPU kernels that unifies template-based refactoring and search-based fine-grained tuning within a single pipeline. By combining hardware resource awareness with measurement feedback, the optimization process is transformed from uncontrolled free-form rewriting into a reproducible, constraint-aware closed loop of parameterized search and validation.
(2) We design a semantics–parameter hierarchical optimization mechanism. At the semantic level, semantics-preserving rewriting improves the structural performance upper bound; at the parameter level, resource-constraint derivation and search-based optimization drive the kernel toward near-optimal mappings, yielding stable performance gains under correctness gating.
(3) We implement an end-to-end closed-loop tuner with multi-agent collaboration, covering semantics-level refactoring, templating, feasible-region derivation, tuning search, correctness verification, and performance measurement. Measurement signals further guide subsequent semantic rewriting, enabling a traceable and reproducible iterative optimization process.
@@ 2 Related Work
Multi-Agent Systems (MAS). MAS consists of multiple interacting agents that collaborate to solve complex shared tasks beyond the capability of a single agent. This paradigm naturally aligns with programming-oriented workflows: complex software development processes can be decomposed into subtasks such as planning, implementation, testing, and profiling, which can then be assigned to different agents for coordinated and iterative progress. In recent years, a growing body of work has explored multi-agent frameworks (e.g., AutoGen, Trace, MetaGPT), demonstrating strong capability and empirical effectiveness on benchmark tasks including mathematical reasoning and code generation [wu2024autogen, cheng2024trace, hong2023metagpt, li2023camel, huang2023agentcoder, qian2024chatdev, wei2024agentinterfaces, yang2024sweagent]. However, systematic application of MAS to parallel code optimization remains relatively limited. A fundamental reason is that performance optimization for modern accelerators relies heavily on specialized knowledge of hardware microarchitecture, parallel execution models, and memory hierarchies, with complex optimization objectives and constraints. As a result, while multi-agent approaches are promising, they also introduce new challenges in modeling and coordination.
@@
Compiler- and learning-based methods for GPU kernel optimization. For decades, parallel code optimization has been driven primarily by compiler frameworks and domain-specific languages (DSLs). Representative systems include Halide, TVM, MLIR, TensorFlow XLA, and NVIDIA CUTLASS [ragankelley2013halide, lattner2020mlir, girija2016tensorflow, thakkar2023cutlass, rotem2018glow, weirichard2017dlvm, jia2019beyonddatamodel, zheng2021neoflow, wei2025taskbasedamr]. These systems provide high-level abstractions for expressing tensor computations and leverage compiler optimization pipelines to perform scheduling transformations, operator fusion, and memory-hierarchy mapping, thereby reducing the cost of manual optimization.
@@
To further approach expert-level performance, autotuning has become a key technique. Frameworks such as AutoTVM, Ansor, AMOS, KernelTuner, and OpenTuner explore large optimization spaces via search and machine learning to identify hardware-specialized configurations [chen2018tvm, zheng2020ansor, zheng2022amos, vanwerkhoven2019kerneltuner, ansel2014opentuner]. Building on this foundation, recent systems (e.g., Triton, Mirage, ThunderKittens) further expand the design space of the “DSL + tuning” paradigm. For instance, Triton introduces a tile-level intermediate representation coupled with autotuning, enabling generated kernels to reach performance close to hand-optimized implementations under certain conditions [tillet2019triton, wu2025mirage, spector2025thunderkittens]. Despite this progress, compiler/DSL-driven approaches still face two prominent limitations. First, without substantial tuning effort, they often struggle to consistently achieve expert-level performance, and their generalization across hardware platforms remains insufficient[spector2025thunderkittens]. Second, these systems typically rely on relatively rigid compilation pipelines, which incur high development and maintenance costs and demand significant engineering investment. In other words, there remains a clear tension between high-performance attainability and system scalability/portability.
@@
LLM-driven high-performance code generation and optimization. Early work on large language model (LLM)-based code generation primarily targeted general-purpose programming tasks (e.g., AlphaCode) and demonstrated promising potential[li2022alphacode]. In recent years, the focus has increasingly shifted toward performance-oriented, domain-specific code generation, spanning vectorization[taneja2025llmvectorizer, zheng2025vectrans], assembly-level optimization[wei2025assemblyrl], parallel program generation for DSLs[wei2025agentinterfaces, wei2024agentinterfaces, wei2025mapple], and tensor program optimization[zhai2024tensorlanguage]. Among these directions, automatic generation of high-performance GPU kernels has emerged as a particularly active line of research[ouyang2025kernelbench, li2025tritonbench]. Unlike general-purpose code generation, performance optimization offers verifiable reward signals (e.g., latency, throughput, and bandwidth utilization). Consequently, an iterative optimization paradigm of “generate–evaluate–improve” has become increasingly prevalent: the model first produces candidate kernels, which are then refined through a feedback loop involving compilation checks, correctness verification, runtime profiling, or self-reflection[brauckmann2025tensorize, wei2024agentinterfaces, chen2025deepseekr1, wang2025geak, wei2025astra]. However, existing LLM- or multi-agent-based tuning processes largely rely on LLM-driven feedback loops and lack an interpretable and controllable mechanism for fine-grained parameter tuning. This often causes key parameter boundaries and interaction effects under hardware resource constraints to be missed, leading to incomplete exploration and high variance in performance.
@@ 3.1 Task Definition
The goal of GPU code tuning is to generate an optimized parameterizable templated kernel and to search for the best general configuration and specialized configuration, so that it runs faster than the original GPU kernel while preserving functional correctness. We formalize the task as follows.
Correctness
Let the input domain be $X$ and the output space be $Y$. The original kernel is
$S:X\rightarrow Y.$
Ideally, the kernel $\tilde{S}$ produced by the two-level optimization should be equivalent to the original kernel over the entire input domain:
$\forall x\in X:\ \tilde{S}(x)=S(x).$
Considering floating-point errors, we allow approximate equivalence under a metric $d$ and a tolerance $\varepsilon\geq 0$:
$\forall x\in X:\ d(\tilde{S}(x),S(x))\leq\varepsilon.$
In practice, correctness is verified on a finite test set $T=\{x_{i}\}_{i=1}^{m}$:
$\max_{1\leq i\leq m}\ d(\tilde{S}(x_{i}),S(x_{i}))\leq\varepsilon.$
Performance Measurement and Evaluation
Let $T_{h}(S,x)$ denote the runtime of kernel $S$ on hardware $h$ for processing input $x$. For each input $x\in T$, we focus on the best achievable performance over the feasible parameter set:
$T_{h}^{*}(x)=\min_{\theta\in\Theta_{\mathrm{feasible}}}\ T_{h}\!\left(\mathrm{Instantiate}(\tilde{S},\theta),\,x\right).$
Here, $\Theta_{\mathrm{feasible}}$ denotes the set of parameter assignments that satisfy both correctness and hardware resource constraints. We define the speedup over the original kernel $S$ as
$\sigma^{*}(x)=\frac{T_{h}(S,x)}{T_{h}^{*}(x)}=\max_{\theta\in\Theta_{\mathrm{feasible}}}\ \frac{T_{h}(S,x)}{T_{h}\!\left(\mathrm{Instantiate}(\tilde{S},\theta),\,x\right)}.$
To aggregate results over the test set $T$, we use the geometric mean speedup:
@@
Here, $m$ denotes the number of samples in the test set $T$. The optimization objective is to maximize this speedup while preserving correctness.
3.2 Multi-agent System
@@
Algorithm 1 presents a two-level GPU tuning pipeline driven by four collaborative agents. The procedure first constructs a finite test set $T$ from the baseline kernel $S_{0}$ and fixes $S_{\mathrm{ref}}=S_{0}$ as the semantic reference. Over $R$ iterations, the planning agent generates semantic-level suggestions $g$ based on measurement feedback, and the generation agent performs semantics-preserving rewriting and templates the kernel into $\tilde{S}_{r}$. The testing agent applies correctness gating under tolerance $\epsilon$ to filter out candidates that violate semantic constraints. For templates that pass the gate, the tuning agent derives a feasible parameter space $\theta_{r}^{\mathrm{feasible}}$ under the hardware resource limits $r(h)$, searches configurations to minimize runtime, and produces a deployable general configuration $\theta_{r}^{\mathrm{gen}}$ and a specialized configuration table $D_{r}^{\mathrm{spec}}$. The system logs results from all rounds and returns the globally optimal $(\tilde{S}^{*},\theta^{\mathrm{gen}*},D^{\mathrm{spec}*})$, enabling traceable and reproducible optimization.
Algorithm 1: A Two-Level GPU Code Tuner: Semantic Refactoring and Search-Based Optimization
Input:
Baseline GPU kernel $S_{0}:\mathcal{X}\rightarrow\mathcal{Y}$
Target hardware $h$ and resource upper-bound vector $r(h)$
Distance metric $d(\cdot,\cdot)$ and tolerance $\epsilon\geq 0$
Searchable parameter vector $\theta\in\Theta$ (discrete candidate space)
Number of semantic rounds $R$
Define:
[…]
$\mathit{Log}\triangleright$ List of $(\textit{round},\,\textit{code},\,\textit{pass},\,\textit{performance},\,\textit{signals})$ for all iterations
Output:
Best template $\tilde{S}^{*}$, best general configuration $\theta^{\mathrm{gen}*}$, best specialized table $D^{\mathrm{spec}*}$, and $\mathit{Log}$
Procedure:
1: $T\leftarrow\texttt{TestingAgent}.\mathrm{GenerateTests}(S_{0})\triangleright$ Initialization: construct a finite test set
2: $S_{\mathrm{ref}}\leftarrow S_{0}\triangleright$ Fix the semantic reference
3: $S_{\mathrm{prev}}\leftarrow S_{0}$; $m_{\mathrm{prev}}\leftarrow\emptyset\triangleright$ Initialize previous code and signals
4: $\mathit{Log}\leftarrow[\,]\triangleright$ Initialize the log
5: $(\tilde{S}^{*},\theta^{\mathrm{gen}*},D^{\mathrm{spec}*})\leftarrow\mathrm{None}$; $score^{*}\leftarrow+\infty\triangleright$ Initialize the incumbent best
[…]
7: Level 1: semantic refactoring
8: if $r>0$ then
9:   $g\leftarrow\texttt{PlanningAgent}.\mathrm{Suggest}(m_{\mathrm{prev}})\triangleright$ Propose semantic-level suggestions
10:   $S_{r}\leftarrow\texttt{GenerationAgent}.\mathrm{Rewrite}(S_{\mathrm{prev}},g)\triangleright$ Semantics-preserving rewriting
[…]
12: end if
13: $pass_{r}\leftarrow\texttt{TestingAgent}.\mathrm{Validate}(S_{r},S_{\mathrm{ref}},T,d,\epsilon)\triangleright$ Correctness gating
[…]
15: end if
16: Level 2: parameter search within the resource-feasible region
17: $\tilde{S}_{r}\leftarrow\texttt{GenerationAgent}.\mathrm{Templatize}(S_{r},\theta)\triangleright$ Produce a parameterizable template
18: $\Theta_{r}^{\mathrm{feasible}}\leftarrow\texttt{TuningAgent}.\mathrm{DeriveFeasibleSpace}(\tilde{S}_{r},h,r(h))\triangleright$ Respect resource limits
19: for each $x\in T$ do
20:   $\theta_{r}^{*}(x)\leftarrow\texttt{TuningAgent}.\mathrm{Search}(\tilde{S}_{r},x,\Theta_{r}^{\mathrm{feasible}})\triangleright$ Minimize runtime under constraints
21: end for
22: $\theta_{r}^{\mathrm{gen}}\leftarrow\texttt{TuningAgent}.\mathrm{SelectGeneral}(\{\theta_{r}^{*}(x)\}_{x\in T},\tilde{S}_{r},T)$
23: $D_{r}^{\mathrm{spec}}\leftarrow\texttt{TuningAgent}.\mathrm{BuildSpecialized}(\{\theta_{r}^{*}(x)\}_{x\in T},T)$
24: $perf_{r}\leftarrow\texttt{TestingAgent}.\mathrm{EvaluatePerformance}(\tilde{S}_{r},\theta_{r}^{\mathrm{gen}},D_{r}^{\mathrm{spec}},S_{\mathrm{ref}},T,h)$
25: $m_{r}\leftarrow\texttt{TestingAgent}.\mathrm{CollectSignals}(\tilde{S}_{r},\theta_{r}^{\mathrm{gen}},D_{r}^{\mathrm{spec}},T,h)$
26: $\mathrm{Append}(\mathit{Log},(r,S_{r},\tilde{S}_{r},\theta_{r}^{\mathrm{gen}},D_{r}^{\mathrm{spec}},pass_{r},perf_{r},m_{r}))$
27: $score_{r}\leftarrow\texttt{PlanningAgent}.\mathrm{Score}(perf_{r})\triangleright$ Negative geometric-mean speedup
28: if $score_{r}<score^{*}$ then
29:   $(\tilde{S}^{*},\theta^{\mathrm{gen}*},D^{\mathrm{spec}*})\leftarrow(\tilde{S}_{r},\theta_{r}^{\mathrm{gen}},D_{r}^{\mathrm{spec}})$; $score^{*}\leftarrow score_{r}$
30: end if
31: $S_{\mathrm{prev}}\leftarrow S_{r}$; $m_{\mathrm{prev}}\leftarrow m_{r}\triangleright$ Update state
32: end for
33: return $(\tilde{S}^{*},\theta^{\mathrm{gen}*},D^{\mathrm{spec}*},\mathit{Log})$